OpenStudy (anonymous):

What should be done to both sides of the equation in order to solve p - 17 = -23? The number 17 should be subtracted. The number 17 should be added. The number -23 should be subtracted. The number -23 should be added.

OpenStudy (anonymous):

any ideas ?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no.... p - 17 = -23 you want p by itself, and if you notice, if you add 17 to both sides you can get p by itself p - 17 + 17 = -23 + 17--combine like terms p = -23 + 17 p = -6 so you would add 17 to both sides understand ?

OpenStudy (anonymous):

no really lol

OpenStudy (anonymous):

so B

OpenStudy (anonymous):

?

OpenStudy (anonymous):

p - 17 = -23 remember this, if you want something moved to the other side of the equal sign, do the opposite of what it is showing. Here you want -17 on the other side. So you would add 17, making the 17's on the left side cancel out. p - 17 + 17 = -23 + 17...and then just combine like terms and solve for p. If your problem said : p + 17 = -23.....you want the 17 on the other side, so do the opposite...the opposite of + is -, so you would subtract 17 from both sides. p + 17 - 17 = -23 - 17...and the 17's cancel out on the left side.

OpenStudy (anonymous):

yes...answer is B

OpenStudy (anonymous):

ok